# Generalizing multiple memories from a single drive: The hysteron latch

### v3

@@ 1.4 Nesting Memories of Asymmetric Driving
Figure 6: Graphical analysis of multiple memories. Each latching pair like Fig. 4D corresponds to a point on a plane, according to its upper thresholds. (A) An infinite ensemble of two-hysteron latches with continuously-distributed thresholds, all initialized to the --- - state. The thresholds are labeled $\gamma_{A}$, $\gamma_{B}$ for convenience. In accordance with Fig. 4D, $\gamma_{B}>\gamma_{A}>0$. (B) “Template” for how an asymmetric cycle with amplitude $\gamma_{0}$ changes the states of latches in the ensemble. (C) The ensemble in (A), after one cycle with amplitude 4% forms a memory. (D) The ensemble encodes two memories after cycles with 4% and 2% amplitude. (E) Readout of the ensemble in (D), reporting the fraction of hysterons in different states, in latches that conform to Fig. 4D and/or Fig. 5A. Curves show readout of all latches (“all”), and of ensembles generated separately to match each labeled region in (D).
@@
The equivalence in Fig. 4 between a single hysteron under symmetric driving and a frustrated pair under asymmetric driving suggests that their multiple-memory capacities may be understood in the same way. Indeed, our method can describe the return-point memory of symmetric cycles that begin with positive strain, via the change of variables $\gamma_{A}\to-\gamma^{-}$, $\gamma_{B}\to\gamma^{+}$. However, our scheme is distinct from earlier graphical analyses of return-point memory for arbitrary driving (neither symmetric nor asymmetric)  (?, ?, ?, ?).
Unlike this idealized picture, in a disordered ensemble of physical hysterons such as an amorphous solid, the distributions of $\gamma_{A}$ and $\gamma_{B}$ obtained will be non-uniform and their range will be finite, according to the particular physics of the system. However, if these distributions are held constant, the ability to encode more memories within the same range—i.e. to form much finer stair-steps than in Fig. 6D—is limited by the number of hysterons in the ensemble. In the Supplementary Text figure S5 we extend our analysis to arbitrarily many memories, and we consider the lengths and areas of the stair-steps of Fig. 6D to show that, as with return-point memory  (?, ?), the maximum number of nested memories scales with the square root of the ensemble size.
Designed systems and sequence recognition — The two-dimensional analysis in Fig. 6 can be made nearly one-dimensional if the hysterons’ parameters can be specified: only some latches near the $\gamma_{A}=\gamma_{B}$ line are needed to encode the stair-step signature of nested memories, and the rest are redundant. We demonstrate this idea by constructing an ensemble. First, we divide the entire range of expected amplitude values into $M$ non-overlapping intervals: $[\tilde{\gamma}_{0},\tilde{\gamma}_{1}),[\tilde{\gamma}_{1},\tilde{\gamma}_{2}),\dots,[\tilde{\gamma}_{M-1},\tilde{\gamma}_{M})$. The two endpoints of each interval then become $\gamma_{A}$ and $\gamma_{B}$ for each of $M$ latches. As in the preceding discussions, finding the $m$th latch latch in the ++-+ - state is evidence that a cycle with an amplitude between $\gamma_{A}=\tilde{\gamma}_{m-1}$ and $\gamma_{B}=\tilde{\gamma}_{m}$ was applied, and that no cycle exceeding $\tilde{\gamma}_{m}$ has been applied since then. Thus $M$ latches digitize and store $M$ distinct amplitudes—a linear scaling, instead of the square-root scaling in a disordered ensemble. Together, the latches’ states are bits that distinguish one sequence of nested amplitudes from among $2^{M}$ possibilities. While a conventional single-dial combination lock does not rely on hysterons (instead using a series of wheels to store different values)  (?, ?), our analysis shows that it is possible to design interacting hysterons so that the functionality of a combination lock is obtained regardless of driving type.
2 Discussion
@@
Together, these two behaviors point to a principle even more generic than return-point memory: robust, nested memories arising from units that “latch” at some input value and reset at a larger value, as in the parallel diagrams of Fig. 4(B, D). For a single hysteron, this pattern is realized by the asymmetric placement of the flipping strains around 0, while for the latching pair, one hysteron cannot return to a “down” state until a large deformation drives another hysteron “up”—the essence of a frustrated interaction.
Recent progress in creating and describing interacting mechanical hysterons  (?, ?, ?, ?, ?) has now led to experiments with highly tunable interactions  (?, ?). Although the interactions in these experimental realizations may be more complex than in Eq. 1, the asymmetric latching behavior comes from a simple relationship among transition thresholds in an interacting pair (Eq. 2) and so lends itself to a variety of experimental systems, as was recently demonstrated with bistable rotors  (?). We have shown that achieving nested memories and sequence recognition in these systems does not require one to fine-tune the competing interactions among multiple hysterons, but instead comes from connecting independent latching pairs to a common drive.
@@
Our work also points to new opportunities for the study of glassy matter. Even though frustrated interactions are essential to the physics of amorphous solids  (?, ?), crumpled sheets  (?), and some magnetic systems  (?, ?, ?) and mechanical metamaterials  (?, ?), in existing memory studies frustration has largely been relegated to perturbing return-point memory. Our results show that a simple change to the driving protocol can suppress return-point memory and reveal a rich, intelligible, and distinctly glassy form of memory. Because this memory arises from a single dominant mechanism, even in larger systems, experiments and molecular dynamics simulations can characterize interaction strengths by tracking individual relaxations while varying the amplitude and origin of asymmetric driving. More generally, the readout method is based on differences, and so we look forward to results like those in Fig. 3B that quantify frustration in macroscopic samples via measurements of magnetization, light scattering, or even image subtraction  (?).
@@
Our study adds to the evidence that frustrated matter can remember what return-point memory must forget—that weakly breaking return-point memory tends to expand memory capacity. This hypothesis is also supported by studies of glassy systems’ ability to divide the frequency of driving  (?, ?, ?, ?) or retain vestiges of erased memories  (?)—although in those examples the mechanisms are unclear or lack a common motif, and they require $N\geq 3$ hysterons. By contrast, the two-hysteron latch is a singular mechanism that is as small as possible, yet scales linearly to store arbitrarily long sequences. The elementary principles and designs emerging from our work and from other recent studies hold promise for new kinds of mechanical information-processing systems as useful, robust, and ubiquitous as the venerable combination lock.
Acknowledgments
We thank Martin van Hecke, Joseph Paulsen, Mahesh Bandi, and Surendra Padamata for insightful conversations.
Funding:
CWL was supported by the US Department of Energy, Office of Science, Basic Energy Sciences, under Grant DE-SC0020972. TRJ was supported in part by a Remote Innovation Grant from the Student Engagement Network at Penn State, and by a grant from the Schreyer Honors College at Penn State.
Author contributions:
All authors wrote the paper. CWL performed theoretical analyses, TRJ and NCK performed simulations and analyzed results, and NCK designed the study.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data needed to evaluate the conclusions in this paper are present in the paper, the Supplementary Materials, and in an archive deposited at
https://doi.org/10.5281/zenodo.14002707.
Supplementary materials
Supplementary Text
Figs. S1 to S5
Table S1
Data S1
Supplementary Materials for
Generalizing multiple memories from a single drive: The hysteron latch
Chloe W. Lindeman, Travis R. Jalowiec, Nathan C. Keim
Corresponding author. Email: keim@psu.edu
This PDF file includes:
Supplementary Text
Figs. S1 to S5
Table S1
Other Supplementary Materials for this manuscript:
Data S1
Supplementary Text
Effect of initial condition
In the main text we consider asymmetric driving with positive strain only, and all hysterons beginning in the “–” state—opposite the polarity of driving, as shown in Fig. 1. Here we describe simulations that begin in the “+” state, as though $\gamma\to\infty$; in these simulations we drive the hysterons to $\gamma=0$ and then write and read memories as in Figs. 1 and 2. Figure S1 shows these results along with the curves from Figs. 3 and 4(A, C). The effect on symmetric driving in Fig. S1(A, C) is subtle. However, the effect of initialization on return-point memory under asymmetric driving in Fig. S1(B, D) can be dramatic. In simulations without interactions ($J_{ij}=0$) the asymmetric curve initialized with the same (positive) sign shows no memory at all—it has a fractional difference of 0 for all readout strains. This is a consequence of return-point memory: because $\gamma$ is bounded between $\infty$ and 0, visiting the turning point at 0 after each cycle restores the system to its state before encoding. By contrast, the second-order memory mechanism illustrated in Fig. 4d is not return-point memory, and its behavior for $\gamma\leq 4\%$ is the same whether we first approach $\gamma=0$ from +$\infty$+ ∞ or $\infty$- ∞. Thus in Fig. S1B, the two asymmetric protocols each form a memory, and their readout results agree for $\gamma_{\text{read}}<4\%$. Differences at larger $\gamma_{\text{read}}$ are due to hysterons that were not affected by writing and readout, and that retain only a memory of the initial condition.
Figure S1: Readout curves showing role of initialization for asymmetric driving. Data are reproduced from Figs. 3 and 4(A,C), with the addition of frustrated ensembles starting from an all-positive initial state (i.e. the same polarity as driving). (A) Readouts after writing a single memory at $\gamma=4\%$. (B) Readouts after writing memories at 4% and 2%. Positive initialization has no effect on the two-hysteron mechanism that gives rise to the signal for $\gamma_{\text{read}}\leq 4\%$, while it changes the slope of the response at larger $\gamma_{\text{read}}$.
Asymmetric driving with same-polarity initialization is thus a special case for return-point memory, and for simplicity we have omitted it from the main text. Other initialization protocols  (?) will generally be able to form memories of single amplitudes under asymmetric driving, even in systems with perfect return-point memory. We note that the readout behavior past the stored memory (here, at $\gamma_{\text{read}}=4\%$) is itself a kind of memory of initial conditions.
System parameters and the interaction strength scale
There are three strain scales that play a role in the effect of interactions on a pair of hysterons: the amount of hysteresis before adding interactions $\gamma^{+}-\gamma^{-}$ (sometimes called the hysteron length), the interaction strength scale $J_{0}$, and the driving amplitude, taken here to be a few percent, $\gamma_{0}\approx 0.04$. The effect of the driving amplitude is to select which hysterons are involved in the writing and readout process: only hysterons with hysteresis $\gamma^{+}-\gamma^{-}\sim\gamma_{0}$ can contribute. Here, the hysteresis is effectively chosen from $[0,0.2]$, and the contributing hysterons are further cut down by $\gamma_{0}\approx 0.04$. A natural non-dimensional scale of interaction strength would therefore be relative to the driving amplitude: $J^{*}\sim J_{0}/\gamma_{0}$. In our scaling study, $J_{0}$ ranges from $10^{-5}$ to $10^{-2}$, giving $J^{*}$ from $\sim 3\times 10^{-4}$ to $\sim 0.3$.
Note that in other cases where hysteresis is small relative to the driving amplitude, the non-dimensional scale of interaction strength might be better defined as the ratio of $J_{0}$ to the hysteron length.
Readout curves for various $J_{0}$ and $N$
Figure S2 shows the readout curves, generated in the manner of Fig. 4c, that were used to measure memory signal strength for Fig. 4f. The “maximum memory signal” plotted in the Fig. 4f refers to the maximum value of each curve for $\gamma_{\text{read}}<4\%$, and is measured only on curves that are non-monotonic, meaning that some larger values of $J_{0}$ are not plotted in the paper. We also omit measurements for $N=2$, $J_{0}<10^{-4}$, for $N=3$, $J_{0}<4.6\times 10^{-5}$, and for $N=9$, $J_{0}<2.2\times 10^{-5}$ because in those cases the entire signal comes from at most 1 hysteron, making a measurement difficult to interpret.
Figure S2: Single-memory readout curves for various interaction strength scales $J_{0}$ and number of mutually-interacting hysterons $N$. Each panel corresponds to a curve in Fig. 4f: (A) $N=2$; (B) $N=3$; (C) $N=9$; (D) the “$N=9\pm$” case where $J_{ij},J_{ji}$ pairs are made positive with probability 1/2. The legend identifies the value of $J_{0}$ for each curve and matches the vertical order of the curves; larger $J_{0}$ generally result in larger fractional differences. Fractional differences of zero are not shown on these logarithmic axes. Each curve was generated from an ensemble of $10^{9}$ groups of $N$ hysterons.
Transition graphs and generalized inequalities
A particular group of hysterons can be represented as a transition graph, in which nodes are states of the hysterons (for example, +++++ + or ++-+ -) and arrows represent transitions between states for increasing and decreasing strain. For two hysterons, all such transition graphs are enumerated in  (?). Of those, only three can contribute to the non-zero readout signal explored in Fig. 4. These transition graphs are shown in Fig. S3, with example values of the flipping strains given for each arrow in the cases observed in simulation. These example flipping strains satisfy Eq. 2.
Figure S3: Two-hysteron transition graphs that give rise to non-zero readout below the remembered strain. Transitions encountered upon increasing (decreasing) strain are in red (blue). Examples of strain values that give rise to such a behavior are shown. (A) Transition graph corresponding to the example in Fig. 4d, also shown in Fig. 4e. (B) A second contributing transition graph, which occurs when $\gamma_{1}^{-}(+)>\gamma_{2}^{+}(+)$, also shown in Fig. 4e. This causes a so-called “horizontal” avalanche when the second hysteron flips to the $+$ state under increasing strain  (?). (C) A third contributing transition graph, which occurs when $\gamma_{2}^{-}(-)>\gamma_{1}^{-}(+)$. This causes a so-called “vertical” avalanche when the first hysteron flips to the -- state under decreasing strain. No example strains are given because strictly negative interactions in the $N=2$ pairs simulated prevented such a graph from contributing to readout. (D) Example hysterons in the strain-space format of Fig. 4(B,D) of the main text and corresponding to the type of transition graph shown in (B) (the flipping strains shown are not the exact values marked in (B); rearrangements have been spaced out for clarity).
Note that, in practice, the transition graph in Fig. S3B did not contribute substantially to the readout signals, as it requires large interaction strength relative to the length $\gamma_{2}^{+}-\gamma_{2}^{-}$ of hysteron 2. We quantify this contribution in the bottom row of Table S1. The transition graph in Fig. S3C did not contribute at all because it requires breaking the rule used throughout the main text that each $J_{ij},J_{ji}$ pair must have the same sign  (?).
$N$ $2$ $3$ $9$ $9\pm$ (a) $9\pm$ (b)
$P(\text{coop})$ 0 0 0 $1/2$ $1/2$
$J_{0}$ $1.00\times 10^{-2}$ $2.00\times 10^{-3}$ $2.00\times 10^{-3}$ $2.00\times 10^{-3}$ $1.00\times 10^{-2}$
$P(\text{sig})$ $2.07\times 10^{-5}$ $2.19\times 10^{-6}$ $8.71\times 10^{-6}$ $4.72\times 10^{-6}$ $7.82\times 10^{-5}$
$P(\text{sig})/(N-1)(1-P(\text{coop}))J_{0}^{2}$ $0.21$ $0.27$ $0.27$ $0.30$ $0.20$
$P(\text{kine}\mid\text{sig})$ $1$ $0.98$ $0.91$ $0.82$ $0.67$
$P(\text{param}\mid\text{kine})$ $1$ $0.24$ $7.21\times 10^{-2}$ $4.57\times 10^{-2}$ $4.69\times 10^{-2}$
$P(\text{coop}\mid\text{kine})$ $1.22\times 10^{-2}$ $6.09\times 10^{-2}$
$\langle J_{12}\rangle/J_{0}$ $-0.78$ $-0.80$ $-0.80$ $-0.79$ $-0.77$
$\langle J_{21}\rangle/J_{0}$ $-0.18$ $-0.20$ $-0.20$ $-0.21$ $-0.20$
$P(\text{avalanche}\mid\text{kine})$ $2.24\times 10^{-2}$ 0 0 0 $2.20\times 10^{-2}$
Table S1: The pair mechanism of Fig. 4d dominates memory of asymmetric driving in larger groups of hysterons. Analysis was performed on ensembles of $10^{9}$ groups of size $N$ (columns); for the two “$9\pm$” ensembles, pairs of $J_{ij},J_{ji}$ were made positive with probability $1/2$. Analysis identified hysterons that contributed to the readout signal (“sig”), that matched the kinematics of Figs. 4d and/or S3D (“kine”), and that had cooperative interactions (“coop”) or horizontal avalanches as in Fig. S3D (“avalanche”). $P(\text{sig})$ is also rescaled according to Fig. 4f. Average interaction values $\langle J_{12}\rangle,\langle J_{21}\rangle$ are from hysteron pairs that match kinematics and have a frustrated interaction; subscripts refer to the corresponding hysteron indices in Fig. 4d. $\langle J_{12}\rangle,\langle J_{21}\rangle$ are remarkably consistent, suggesting that memory kinematics of a pair are driven primarily by their interaction, even in larger groups.
To generalize the inequalities in Eq. 2 to include all three transition graphs, we allow for non-frustrated interactions. For the latching mechanism necessary for non-zero readout to occur, we still need one frustrated interaction (without loss of generality, $J_{12}<0$), so the only new case allowed by this relaxation of strictly frustrated interactions is $J_{21}>0$.
The general picture that emerges from the three contributing diagrams (shown in Fig. S3A–C) is (1) that the hysterons must travel through all four possible states in a cycle of sufficient size (rather than, for example, going through ++-+ - on both the way up and the way down); and (2) that the transition back from ++-+ - to --- - is outside the range of driving, that is, below zero. This sets up the following inequalities:
  • If we define hysteron 1 as the hysteron that flips forward first, then hysteron 1 should also be the first to flip back: $\gamma^{-}_{1}(+)>\gamma^{-}_{2}(+)$. This makes it possible to travel through all four states in a shear cycle.
  • The transitions from +++++ + to +-+- + and from +-+- + to --- - must occur at $\gamma>0$. In other words, since we have already determined that hysteron 1 is the first to flip back to --, $\gamma^{-}_{1}(+)>0$ and $\gamma^{-}_{2}(-)>0$.
  • The transition from ++-+ - to --- - must be below zero: $\gamma^{-}_{1}(-)<0$. This prevents the pair from always returning to the same state when the strain is returned to 0.
More concisely, the necessary and sufficient set of inequalities to get non-zero readout below the remembered strain for two interacting hysterons that have been driven asymmetrically is
$\gamma^{-}_{1}(-)$ $<0<\gamma^{-}_{1}(+)$ (S1)
$\gamma^{-}_{2}(+)$ $<\gamma^{-}_{1}(+)$ (S2)
00 0 $<\gamma^{-}_{2}(-),$ (S3)
which is satisfied by graphs S3(A–C) and in Figs. 4d and S3D, and of which Eq. 2 is the special case for $J_{21}<0$.
Frustrated-Pair Mechanism in Larger Systems
The scaling analysis in Fig. 4f suggested that the memory of asymmetric driving in larger groups of $N$ mutually-interacting hysterons (Fig. S2) was dominated by the action of frustrated pairs. While for $N=2$ one can predict memory behavior by inspecting the model parameters (Eqs. 1 and 2), for $N>2$ the numerous interactions and metastable states make this impractical. Instead, we can investigate the origins of memory readout signals in larger groups and ask whether the mechanism for memory in a pair is at work, by examining kinematics: the transitions of individual hysterons within larger groups, as a memory is written and read with the asymmetric protocol in Fig. 2c.
Table S1 summarizes our analysis of the kinematics in ensembles of $10^{9}$ groups, generated according to the same schemes as the inset of Fig. 3b and Fig. S2. We fix the maximum interaction strength at $J_{0}=2\times 10^{-3}$, except $J_{0}=10^{-2}$ for $N=2$, the $N$ for which the memory behavior is least likely in a group. We first drive an ensemble with the single-memory asymmetric protocol in Fig. 2a. We identify the probability $P(\text{sig})$ that a given hysteron contributes to a non-monotonic memory readout signal, as in Fig. 4c — specifically, that its state after the cycle with $\gamma_{\text{read}}=4\%$ matches its state at the beginning of readout, and that it is in a different state at the end of at least one intervening cycle.
Next, each hysteron that counts toward $P(\text{sig})$ belongs to several interacting pairs (for $N$ hysterons, $N-1$ pairs). For each pair, we extract the sequence of states that the pair visits in each cycle of readout. Figure 4b establishes that the memory behavior of interest arises from the distinct sequences obtained during small-, medium-, and large- amplitude cycles. To check whether these sequences match the progressions in Fig. 4d or Fig. S3D, our analysis implements the finite-state machine in Fig. S4; this also checks for a (rare) variant made possible by the avalanche behavior. We then ask: of the hysterons that contribute to the readout signal, what fraction $P(\text{kine}\mid\text{sig})$ belongs to exactly one pair that matches these kinematics? Remarkably, in Table S1 we report that $P(\text{kine}\mid\text{sig})\sim 1$ even in large groups with cooperative and frustrated interactions (“$9\pm$”).
Figure S4: Testing the kinematics of a hysteron pair. A finite-state machine tests whether the kinematics of a memory-encoding pair during readout match the readout sequence shown in Fig. 4D or Fig. S3D. Each arrow represents one cycle of asymmetric driving and shows the sequence of states (if any) during that cycle; each node represents the state at the beginning/end of the cycle. Readout begins in the --- - state at left and proceeds with increasing $\gamma_{\text{read}}$. To be accepted as a kinematics match, a hysteron pair must follow only an allowed sequence of states, and must be in the right-hand --- - state after the cycle with $\gamma_{\text{read}}=4\%$ (the strain amplitude that was written). The longer sequences from ++-+ - to --- - and from --- - to itself correspond to the uppermost transitions in Fig. S3B.
We can further ask whether the actual mechanism of Fig. 4d, which is based on an isolated pair of hysterons, has any explanatory value for the pairs embedded in larger groups. At one extreme, one might hypothesize that these pairs are nearly isolated from their companions, so that their randomly-generated interaction and threshold parameters are often consistent with Eq. 2—a probability we measure as $P(\text{param}\mid\text{kine})$. Instead, Table S1 shows that $P(\text{param}\mid\text{kine})\ll 1$ in large groups, meaning that the effects of other hysterons on the thresholds in Eq. 2 are significant; the behavior cannot be predicted from the two hysterons alone.
At the other extreme, one might hypothesize that a pair of hysterons with memory kinematics is merely a “puppet” of other hysterons with unknown memory behaviors, and that the interaction within the pair is so unimportant that it need not even be frustrated. In the “$9\pm$” ensembles where cooperative interactions occur with probability 0.5, we find that this “puppet” scenario is not supported, since $P(\text{coop}\mid\text{kine})\ll 1$.
Most remarkably, Table S1 shows that for hysteron pairs that match kinematics and have a frustrated interaction, the mean interaction values $\langle J_{12}\rangle,\langle J_{21}\rangle$ are nearly the same fraction of $J_{0}$ in all ensembles. In other words, the kinematics of a memory-forming pair within a larger group are driven by the interactions within that pair, in the same way as for $N=2$. Together, the results in Table S1 strongly support our interpretation of the $J_{0}^{2}$ scaling results in Fig. 4f: the mechanism in Fig. 4d is the dominant way that hysterons remember asymmetric driving, even in large systems with a mixture of cooperative and frustrated interactions.
Generalizing multiple-memory capacity in disordered ensembles
Here we generalize the template method outlined in Fig. 5 of the main text beyond the one- and two-amplitude writing sequences shown there. Figures S5(A–D) reproduce the initial case, the general template, and the one- and two-amplitude cases from Fig. 5(A–D). Figures S5(E, F) show the extension of these cases: it is possible to distinguish histories even when a larger number of amplitudes are written.
Figure S5: Graphical analysis of multiple memories. Each pair of hysterons like Fig. 4D can be represented as a point on a plane, according to its upper thresholds. (A) An infinite ensemble of such pairs with continuously-distributed thresholds, all initialized to the --- - state. The thresholds are labeled $\gamma_{A}$, $\gamma_{B}$ for convenience. In accordance with Fig. 4d, $0<\gamma_{A}<\gamma_{B}$. (B) “Template” for how an asymmetric cycle with amplitude $\gamma_{0}$ changes the states of pairs in the ensemble. (V) The ensemble in (A), after one cycle with amplitude 4% forms a memory. (D) The ensemble encodes two memories after cycles with 4% and 2% amplitude. (E) Ensemble after cycles with 4%, 2%, and 1%. (F) Ensemble after cycles with 4% and 1%. Comparing with (E) reveals the set of pairs (outlined with heavy dashes) that encode the difference between these histories.
Following the argument of Keim and Medina  (?), we consider briefly the memory capacity of a finite, disordered ensemble of frustrated pairs. Figure S5E adds a third memory, and illustrates that nested memories form a stair-step pattern on the plot. If the density of hysteron pairs on the plane is finite, then these stair-steps can be subdivided only so many times before they become indistinct. More precisely, if we generalize our memory values as the descending sequence $\gamma_{n},\gamma_{n-1},\dots,\gamma_{2},\gamma_{1}$, then memory $0<m<n$ from within the sequence will be encoded by the pairs in the rectangle $\gamma_{m}<\gamma_{B}\leq\gamma_{m+1}$, $\gamma_{m-1}<\gamma_{A}\leq\gamma_{m}$. This can be seen by comparing Fig. S5E with the outcome if the 2% memory had not been written (Fig. S5F); the highlighted rectangle encodes this difference. If there is no pair of hysterons within such a rectangle, then the ensemble’s state will be the same whether $\gamma_{m}$ was written or not—the memory capacity has been exceeded. Thus, in a disordered ensemble of $N$ latching pairs, the maximum number of distinct memories $M$ scales as the maximum number of distinct rectangles that can be drawn in this way, such that each rectangle contains at least one pair. The linear spacing of these finely-drawn rectangles then scales as $1/M$, and the area of each one scales as $1/M^{2}$.
In the case of a disordered ensemble with roughly continuous distributions of $\gamma_{A}$ and $\gamma_{B}$, the area of the smallest rectangle that contains one latching pair scales inversely with the number density of the pairs, leading to $1/M^{2}\sim N$, for a memory capacity $M\sim\sqrt{N}$ as in the Preisach model  (?, ?). The presence or absence of each memory corresponds to a different history, and so readout lets one distinguish among $\sim 2^{M}$ possible histories of asymmetric driving.
Caption for Data S1.
Supporting data files. The data archive referenced in the main text, deposited on the Zenodo service and available at https://doi.org/10.5281/zenodo.14002707, contains a directory for each data plot and table in the text and Supplementary Materials. Each directory contains one or more CSV-formatted files containing the values for each curve in that plot. The directory names are listed below, along with notes where applicable:
  • 2b-readout-symm-noninteracting — For clarity we avoided plotting curves that were identical (within random error), as is expected from the perfect return-point memory of these non-interacting hysteron systems. Thus, in both figures, the ’2,4’ curve was not plotted because it cannot differ significantly from the ’4’ curve. We provide the data files ’24.csv’ so that a reader can verify this agreement predicted by theory. In Fig. 2d, the ’4,2’ curve also matched the ’4’ curve, and so it was not plotted, but we provide the data file ’42.csv’ for verification.
  • 2d-readout-asymm-noninteracting — See previous note.
  • 3a-readout-symm-2mem
  • 3b-readout-asymm-2mem
  • 4a-readout-symm-1mem
  • 4c-readout-asymm-1mem
  • 4f-scaling — Data files are the un-rescaled maximum fractions different during readout; to rescale and obtain the values shown on the plot, see caption of Fig. 4f.
  • 5e-readout-segmented
  • S1a-readout-symm-1mem-polarity
  • S1b-readout-asymm-1mem-polarity
  • S1c-readout-symm-2mem-polarity
  • S1d-readout-asymm-2mem-polarity
  • S2a-int_strength_N=2
  • S2b-int_strength_N=3
  • S2c-int_strength_N=9
  • S2d-int_strength_N=9pm
  • table-S1-kinematics

### v4

@@ II.4 Nesting Memories of Asymmetric Driving
The equivalence in Fig. 4 between a single hysteron under symmetric driving and a frustrated pair under asymmetric driving suggests that their multiple-memory capacities may be understood in the same way. Indeed, our method can describe the return-point memory of symmetric cycles that begin with positive strain, via the change of variables $\gamma_{A}\to-\gamma^{-}$, $\gamma_{B}\to\gamma^{+}$. However, our scheme is distinct from earlier graphical analyses of return-point memory for arbitrary driving (neither symmetric nor asymmetric) [3, 7, 23, 5]. In the Supplemental Materials [21] we extend our analysis to arbitrarily many memories and we consider the areas of the stair-steps of Fig. 5d to show that, as with return-point memory, the maximum number of nested memories scales with the square root of the ensemble size.
Designed systems and sequence recognition — The two-dimensional analysis in Fig. 5 can be made nearly one-dimensional if the hysterons’ parameters can be specified: only some latches near the $\gamma_{A}=\gamma_{B}$ line are needed to encode the stair-step signature of nested memories, and the rest are redundant. We demonstrate this idea by constructing an ensemble. First, we divide the entire range of expected amplitude values into $M$ non-overlapping intervals: $[\tilde{\gamma}_{0},\tilde{\gamma}_{1}),[\tilde{\gamma}_{1},\tilde{\gamma}_{2}),\dots,[\tilde{\gamma}_{M-1},\tilde{\gamma}_{M})$. The two endpoints of each interval then become $\gamma_{A}$ and $\gamma_{B}$ for each of $M$ latches. As in the preceding discussions, finding the $m$th latch latch in the ++-+ - state is evidence that a cycle with an amplitude between $\gamma_{A}=\tilde{\gamma}_{m-1}$ and $\gamma_{B}=\tilde{\gamma}_{m}$ was applied, and that no cycle exceeding $\tilde{\gamma}_{m}$ has been applied since then. Thus $M$ latches digitize and store $M$ distinct amplitudes—a linear scaling, instead of the square-root scaling in a disordered ensemble. Together, the latches’ states are bits that distinguish one sequence of nested amplitudes from among $2^{M}$ possibilities. Testing the states of these bits forms the basis for a combination lock.
III Discussion
@@
Together, these two behaviors point to a principle even more generic than return-point memory: robust, nested memories arising from units that “latch“ at some input value and reset at a larger value, as in the parallel diagrams of Fig. 4(b, d). For a single hysteron, this pattern is realized by the asymmetric placement of the flipping strains around 0, while for the latching pair, one hysteron cannot return to a “down” state until a large deformation drives another hysteron “up”—the essence of a frustrated interaction.
Recent progress in creating and describing interacting mechanical hysterons [11, 10, 28, 19, 30] has already led to experiments with highly tunable interactions [19]. Because the two-hysteron latching behavior is described by a simple rule (Eq. 2), it may be implemented in such experiments as a mechanism for multiple memories or sequence recognition.
@@
Our work also points to new opportunities for the study of glassy matter. Even though frustrated interactions are essential to the physics of amorphous solids [31, 14], crumpled sheets [9], and some magnetic systems [24, 25, 32] and mechanical metamaterials [16, 28], in existing memory studies frustration has largely been relegated to perturbing return-point memory. Our results show that a simple change to the driving protocol can suppress return-point memory and reveal a rich, intelligible, and distinctly glassy form of memory. Because this memory arises from a single dominant mechanism, even in larger systems, experiments and molecular dynamics simulations can characterize interaction strengths by tracking individual relaxations while varying the amplitude and origin of asymmetric driving. More generally, the readout method is based on differences, and so we look forward to results like those in Fig. 3b that quantify frustration in macroscopic samples via measurements of magnetization, light scattering, or even image subtraction [23].
@@
Our study adds to the evidence that frustrated matter can remember what return-point memory must forget—that weakly breaking return-point memory tends to expand memory capacity. This hypothesis is also supported by studies of glassy systems’ ability to count repeated cycles [20, 29, 26, 18] and retain vestiges of erased memories [29]—although in those examples the mechanisms are unclear or lack a common motif, and they require $N\geq 3$ hysterons. By contrast, the two-hysteron latch is a singular mechanism that is as small as possible, yet scales linearly to store arbitrarily long sequences. The elementary principles and designs emerging from our work and from other recent studies hold promise for new kinds of mechanical information-processing systems as useful, robust, and ubiquitous as the venerable combination lock.
Acknowledgements.
We thank Martin van Hecke, Mahesh Bandi, and Surendra Padamata for insightful conversations. CWL was supported by the US Department of Energy, Office of Science, Basic Energy Sciences, under Grant DE-SC0020972. TRJ was supported in part by a Remote Innovation Grant from the Student Engagement Network at Penn State, and by a grant from the Schreyer Honors College at Penn State. All authors wrote the paper. CWL performed theoretical analyses, TRJ and NCK performed simulations and analyzed results, and NCK designed the study. The authors declare that they have no competing interests. All data needed to evaluate the conclusions in this paper are present in the paper, the Supplementary Materials, and in an archive deposited at https://zenodo.org/XXXXXX.